# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v3

@@ Abstract
TIPO (Text-to-Image Prompt Optimization) introduces an efficient approach for automatic prompt refinement in text-to-image (T2I) generation. Starting from simple user prompts, TIPO leverages a lightweight pre-trained model to expand these prompts into richer, detailed versions. Conceptually, TIPO samples refined prompts from a targeted sub-distribution within the broader semantic space, preserving the original intent while significantly improving visual quality, coherence, and detail. Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks.
We provide visual results, human preference report to investigate TIPO’s effectiveness. Experimental evaluations on benchmark datasets demonstrate substantial improvements in aesthetic quality, significant reduction of visual artifacts, and enhanced alignment with target distributions along with significant human preference proficiency. These results highlight the importance of targeted prompt engineering in text-to-image tasks and indicate broader opportunities for automated prompt refinement.
@@ 1 Introduction
The rapid proliferation of Text-to-Image (T2I) generative models have revolutionized artistic creation [48, 7, 20, 57, 53, 54, 60, 55, 50, 59, 15, 14, 35, 21, 8]. These models offer direct control over generative visual content via text prompts. To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout. However, the increasing complexity of prompts also forces users to iteratively refine their inputs, making high-quality T2I artwork accessible only to those with significant prompt engineering expertise.
@@
Recently, efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality. A straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner [42]. Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models. This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts. A more effective approach is to train LLMs directly on prompt data collected from model users [5, 17]. While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs. Recent work [26] trains LLM with reinforcement learning, where aesthetic scores of generated images serve as rewards. However, the reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
@@
Figure 1: Comparison of prompt optimization methods using LLM. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
Deviating from existing work, we argue that the optimal prompt lies in T2I models’ training set. Precisely, when a prompt aligns with the data distribution used for model training, the model can better interpret the prompt, resulting in improved text alignment and higher image quality. Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. At its core lies a lightweight language model trained on multiple meticulously designed pretext tasks that transform prompts into tags or sentences from coarse to fine. Leveraging the trained model, TIPO can progressively and seamlessly adapt to various user input types, generating accurate and robust prompts for various T2I models. In addition, our training data is derived from a curated text description dataset comprising over 30 million text-image pair samples and more than 20 billion tokens, ensuring comprehensive coverage of T2I training sets. Extensive experiments on test prompts from in- and out-of-domain demonstrate that TIPO consistently outperforms state-of-the-art methods in fidelity, aesthetics, and text alignment. Detailed analysis shows that TIPO has up to 62.80% win rate in user preference and 25% improvement in runtime efficiency over the second-best baseline. To summarize, our contributions are at least three-fold:
@@
    We introduce TIPO, a prompt optimization framework that leverages the large-scale text data distributions used in text-to-image (T2I) training.
@@
    We train a multi-task language model, which adaptively and progressively refines user inputs from either tag-based or natural language prompts into unified formats, enhancing compatibility across a broad spectrum of T2I models.
@@
    Extensive experiments demonstrate that TIPO achieves superior image quality, strong text alignment, higher human preference, and improved runtime efficiency across both in-domain and out-of-domain prompts, highlighting its great potential in real-world applications.
@@ 2 Related Work
Prompt optimization for T2I models typically leverages language models to refine user inputs. They can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that aim to improve prompt quality across a variety of T2I models.
2.1 Model-specific Strategies
T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference. These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model. For instance, Promptist [26] fine-tunes a pre-trained language model by using CLIP relevance scores as rewards. Similarly, PAE [44] extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning. However, these methods are computationally intensive, often struggling with a larger number of training prompts. Moreover, a model optimized for one specific T2I system may not generalize well to others. In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
2.2 Universal Strategies
To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models. For example, CogView3 [71] employ GLM-4 [24], and Lee et al. [33] employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement. However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models. Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs [5, 63, 17]. Such methods, however, are limited by the inconsistent expertise of users. Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
@@ 3 Preliminaries
Notations. We present the notations in our work:
  • $\mathcal{P}$: The space of all possible prompts.
  • $\mathcal{I}$: The space of all possible images.
  • $p$: User-provided prompt.
  • $p_{o}$: Optimized prompt.
  • $p_{ps}$: Sampled modified prompt derived from p.
  • $\mathcal{M}$: Metadata (e.g., style tags) aiding prompt transformation.
  • $\mathcal{T}$: Transformation function for pre-sampling.(e.g., Random augmentation, TIPO or LLM)
  • $T_{p},T_{ps}$: Sets of tags for original and pre-sampled prompts.
  • $S_{p},S_{ps}$: Ordered lists of natural language descriptions for original and pre-sampled prompts.
Text-to-Image Model. A text-to-image (T2I) model is a mapping function
$f(p)\;\;\to\;\;\mathcal{I}_{p}\subseteq\mathcal{I},$
which maps a user-provided text prompt $p$ to a subset of images $\mathcal{I}_{p}$ corresponding to the prompt.
Problem statement. Let $\mathcal{I}_{u}$ denote the user’s intended distribution over images. The task of Prompt optimization is to generate a prompt $p_{o}$ that minimizes a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and $\mathcal{I}_{u}$:
@@
where $d$ is a suitable metric (e.g., Frechet Inception Distance) comparing two image distributions.
@@ 4.1 Text Set Preparation
Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions. Tag-based captions, such as those in the Danbooru2023 dataset [46, 68], use comma-separated, succinct terms to describe image content. In contrast, NL-based captions, typically generated by language models with visual capabilities [37, 1, 2, 34, 6, 16, 67, 18], may comprise multiple sentences. We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
While the original image captions are fine-grained and detailed, which can yield high-quality images when all elements are used, they often result in prompts that are excessively lengthy or overloaded with information. Such prompts diverge from typical user input and pose alignment challenges. To mitigate this, we construct a simpler subset $T_{s}\subset T$ by removing some tags and sentences from the original set [^1] For long sentences, off-the-shelf text splitters (e.g., FastText) are used to segment them into multiple sentences., as detailed in Section 4.2.
@@ 4.2 Formatted Prompt Construction
We aim to construct prompts in a unified format compatible with existing image captions. First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>. These metadata categories primarily include artist, copyright, aspect ratio, quality, and year (e.g., quality: masterpiece, artist: Picasso). This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
Next, we construct both tag-based and NL-based prompts using text sets $T$. Our design generates both simple (incomplete) and complete prompts for each image, and we train an auto-regressive language model to extend the simple prompts into complete versions. For tag-based prompts, since the tags are largely order-insensitive (i.e., the order has minimal impact on T2I outcomes), we propose a prefix-based dropout strategy. We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption. Then, we construct a simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags. The prompts are constructed as:
@@
Because $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand the simple tag-based prompt into its complete version.
For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information [25] and the order of sentences significantly influences the caption’s semantics. Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order. Let:
@@ 4.3 Text Pre-sampling
We aim to align user input with the constructed high-quality training prompts $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling” - a process that first constrains the target distribution in text space before proceeding to image generation. Conceptually, while pre-sampling strategies generally add more text tokens to user input, the outcomes of different strategies may still vary (as illustrated in Figure 2).
To enrich the image content while mitigating undesired output, we perform pre-sampling based on a language model trained on $p_{o}$. A straightforward pre-sampling strategy is to do text completion, i.e., directly append output tokens to the end of user input. However, we posit that the user input is unlikely to be a part of a good prompt. The na’́ive text-completion strategy prioritizes the user’s original input, resulting in worse outcomes as evidenced by Section 5.6.
Figure 3: TIPO’s task flow within a single generation. Starting from any node, each arrow represents a sequential extension step, with other prompts used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
We propose the core technique of TIPO, a flexible pre-sampling mechanism that divides the prompt optimization process into three interconnected subtasks: enriching tag sequences, extending NL prompts, and refining NL prompts. As shown in Figure 3, TIPO uses both tag and NL prompts interchangeably as input sources to generate detailed outputs in various forms. For instance, TIPO can begin with a short NL prompt to produce a detailed tag sequence, denoted as short_to_tag. Training data for the language model is constructed by concatenating different types of $p_{o}$ with task-specific metadata embedded into the text.
The individual tasks are defined as follows:
  • tag_to_long: Use tags as metadata to generate a new NL prompt.
  • long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  • short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
@@
  • short_to_long: Use a user-provided NL prompt as metadata to generate a refined, detailed prompt.
We also consider composite tasks that can be performed within a single forward pass. These tasks offer the language model a holistic view of the objective while reducing computational overhead:
  • short_to_tag_to_long: Use a user-provided NL prompt or tag sequence as metadata to generate a refined, detailed prompt.
  • short_to_long_to_tag: Use a user-provided NL prompt or generated NL prompt as metadata to extend a tag sequence.
@@
  • tag_to_short_to_long: Use user-provided tags or NL prompts as metadata to generate a refined NL prompt.
@@
We randomly select from the aforementioned tasks during training to enhance model generalization. By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences. Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence:
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\texttt{metadata}=\emptyset\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, \ldots}\right\}\end{array}$
@@
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
4.4 Implementation Details
TIPO can utilize various types of causal autoregressive language models by design. In practice, we leverage the LLaMA architecture [64, 65, 3] for simplicity, with 200M and 500M parameters.
  • TIPO-200M: Pretrained on the Danbooru2023 [46, 69] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.

### v4

@@ Abstract
TIPO (Text-to-Image Prompt Optimization) introduces an efficient approach for automatic prompt refinement in text-to-image (T2I) generation. Starting from simple user prompts, TIPO leverages a lightweight pre-trained model to expand these prompts into richer, detailed versions. Conceptually, TIPO samples refined prompts from a targeted sub-distribution within the broader semantic space, preserving the original intent while significantly improving visual quality, coherence, and detail. Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks. Extensive experiments across multiple domains demonstrate that TIPO delivers stronger text alignment, reduced visual artifacts, and consistently higher human preference rates, while maintaining competitive aesthetic quality. These results highlight the effectiveness of distribution-aligned prompt engineering and point toward broader opportunities for scalable, automated refinement in text-to-image generation.
National Tsing Hua University  Nanyang Technological University  Yonsei University
Onoma AI  Anhui Medical University  Seoul National University  Comfy Org Research
Corresponding Author  Equal Contribution
@@ 1 Introduction
The rapid proliferation of Text-to-Image (T2I) generative models has revolutionized artistic creation (Ossa et al., 2024; Betker et al., 2023; Esser et al., 2024a; Saharia et al., 2022; Ramesh et al., 2021; 2022; Shi et al., 2020; Rombach et al., 2022a; Podell et al., 2024; Sauer et al., 2024; Chen et al., 2024b; a; Li et al., 2024b; Esser et al., 2024b; black-forest-labs, 2024). These models offer direct control over generative visual content via text prompts. To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout. However, the increasing complexity of prompts often forces users to iteratively refine them to convey intent. Moreover, most state-of-the-art T2I models are aesthetically fine-tuned (e.g., on LAION-aesthetics (Podell et al., 2024)) to favor nuanced artistic and stylistic cues, making high-quality T2I artwork mainly accessible to those with significant artistic expertise.
@@
Extensive efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality. As shown in Figure 1(a), a straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner (mañas2024improvingtexttoimageconsistencyautomatic). Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models. This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts. Figure 1(b) shows a more effective approach: training LLMs directly on prompt data collected from model users (AUTOMATIC, 2022; daspartho, 2022). While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs. Recent work (Hao et al., 2023) trains LLMs with reinforcement learning, where aesthetic scores of generated images serve as rewards, as depicted in Figure 1(c). However, reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
@@
Figure 1: Comparison of prompt optimization methods using LLM. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not the T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and diverse prompt generation that fits the target T2I model.
In contrast to prior work, we argue that good prompts should align with the large-scale text distributions of T2I models’ training, including those emphasized in aesthetic fine-tuning. Such alignment allows models to better interpret user intent and leverage their learned priors, resulting in stronger text alignment and overall image quality. Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), a framework that brings prompt optimization into the domain of large-scale multi-task pre-training. TIPO is supported by a curated 30M-pair, 40B-token caption corpus, which is filtered and balanced to maximize compatibility with leading T2I models and to preserve aesthetic quality. On top of this corpus, we design a suite of pretext tasks that reformulate raw user inputs, including both concise natural sentences and tag-based prompts, into enriched and distribution-consistent forms. Through this multi-task sampling pipeline, a lightweight language model expands (rather than fully rewrites) user prompts, preserving original semantics while enriching details that are diverse, edit-friendly, and aligned with T2I training distributions. Extensive experiments on both in-domain and out-of-domain prompts show that TIPO consistently outperforms strong baselines, achieving a 62.8% win rate in human preference (validated by more than 1,400 pairwise comparisons from 221 volunteers), and providing up to a 29.4% runtime efficiency improvement. Figure 1 illustrates the conceptual differences between existing methods and TIPO. To summarize, our contributions are at least threefold:
@@
    We introduce TIPO, a prompt optimization framework that leverages the large-scale text distributions used in text-to-image (T2I) training.
@@
    We train a lightweight multi-task language model that progressively refines both tag-based and natural language inputs into unified prompts, enhancing compatibility across a broad spectrum of T2I models.
@@
    Extensive experiments demonstrate that TIPO achieves superior image quality, stronger text alignment, higher human preference, competitive aesthetic quality, and improved runtime efficiency against strong baselines with SOTA T2I models, highlighting its practical value.
@@ 2 Related Work
Prompt optimization for T2I models typically leverages language models, which can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that improve prompt quality across a variety of T2I models.
Model-specific Strategies
T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference. These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model. For instance, Promptist (Hao et al., 2023) fine-tunes a pre-trained language model by using CLIP relevance scores as rewards. Similarly, PAE (Mo et al., 2024) extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning. However, these methods are computationally intensive, often struggling with a larger number of training prompts. Moreover, a model optimized for one specific T2I system may not generalize well to others. In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
Universal Strategies
To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models. For example, CogView3 (Zheng et al., 2024) employ GLM-4 (GLM et al., 2024), and Lee et al. (2024) employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement. However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models. Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs (AUTOMATIC, 2022; succintly, 2022; daspartho, 2022). Such methods, however, are limited by the inconsistent expertise of users. More recently,  He et al. (2025); Liu et al. (2024b) leverage vision language models to optimize prompts in an iterative loop: a user prompt generates images via a T2I model, the prompt–image pairs are evaluated by the VLM to suggest refinements, and the revised prompts are reapplied to T2I generation for continual improvement. While such iterative refinement can improve quality, the distribution mismatch between VLMs and T2I training data persists. Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
@@ 3 Preliminaries
We present the formal definition of T2I models and the problem statement of this work.
Text-to-Image Model. A text-to-image (T2I) model defines a conditional distribution
$\mathcal{I}_{p}=P(x\mid p),\quad x\in\mathcal{X},$
which maps a prompt $p$ to a distribution of images over the space of all possible images $\mathcal{X}$.
Problem statement. Let $\mathcal{I}_{u}$ denote the user’s intended distribution over $\mathcal{X}$. The task of prompt optimization is to find an optimized prompt $p_{o}$ from the space of all possible prompts $\mathcal{P}$ to minimize a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and the intended distribution $\mathcal{I}_{u}$:
@@
where $d(\cdot,\cdot)$ is a distance between image distributions (e.g., Fréchet Inception Distance).
4 Methodology
@@ 4.1 Text Set Preparation
Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions. Tag-based captions, such as those in the Danbooru2023 dataset (nyanko202, 2023; Yeh, 2024b), use comma-separated, succinct terms to describe image content. In contrast, NL-based captions, typically generated by language models with visual capabilities (Liu et al., 2023; Agrawal et al., 2024; OpenAI, 2024; Li et al., 2024a; Bai et al., 2023; Dai et al., 2024; Xiao et al., 2024; Deitke et al., 2024), may comprise multiple sentences. We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
While the original image captions are fine-grained and detailed, which can yield high-quality images when all elements are used, they often result in prompts that are excessively lengthy or overloaded with information. Such prompts diverge from typical user input and pose alignment challenges. To mitigate this, we construct a simpler subset by removing some tags and sentences from the original prompt 1set, as detailed in Section 4.2.
@@ 4.2 Formatted Prompt Construction
We aim to construct prompts in a unified format compatible with existing image captions. First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>. These metadata categories primarily include style, aspect ratio, quality, and year (e.g., quality: masterpiece, style: Impressionist). This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
Next, we construct both tag-based and NL-based prompts using text sets $T$. Our design generates both simple (incomplete) and complete prompts for each image, and we train an auto-regressive language model to extend the simple prompts into complete versions. For tag-based prompts, since the tags are largely order-insensitive (i.e., the order has minimal impact on T2I outcomes), we propose a prefix-based dropout strategy. We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption. Then, we construct a simpler tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags. The prompts are constructed as:
@@
Here, $p_{s}$ and $p_{o}$ denote the simple and original prompts, respectively. By ensuring that $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand simple tag-based prompts into complete versions.
For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information (Godbole et al., 2024) and the order of sentences significantly influences the caption’s semantics. Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order. Let:
@@ 4.3 Text Pre-sampling
We aim to reformulate user input into forms that better align with the high-quality training text distribution $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling.” A naïve strategy is plain text completion, where tokens are directly appended to the user input in an unstructured manner. Such completion often mirrors the low-quality phrasing of the input, deviates from the distribution of high-quality prompts, and produces inferior generations. On the other hand, a full rewrite risks deviating from the user’s original intention. To balance these issues, TIPO preserves the original input and appends a structured, distribution-consistent expansion. This appended segment is typically paragraph-like or tag-like, making it both informative and easy to edit or remove (see Figure 2 for an illustration and Appendix E for concrete examples).
We propose the core technique of TIPO, a flexible pre-sampling mechanism that decomposes prompt optimization into three subtasks: enriching tag sequences, extending natural language (NL) prompts, and refining NL prompts. For example, a short NL prompt can be expanded into a detailed tag sequence (short_to_tag), as illustrated in Figure 3. We further distinguish between basic tasks, which perform a single transformation, and composite tasks, which chain multiple transformations within a single forward pass. The latter expose the model to more holistic training signals while reducing computational overhead. Table 1 summarizes all tasks and their input–output forms.
Task Description
Basic tasks
tag_to_long Generate a new NL prompt given tags.
long_to_tag Extend a tag sequence given an NL prompt.
short_to_tag Extend a tag sequence given a short/simple prompt $p_{s}$.
@@
short_to_long Generate a refined, detailed NL prompt given a user-provided NL prompt.
Composite tasks
short_to_tag_to_long From a short NL prompt or tags, produce a refined NL prompt.
short_to_long_to_tag From a short or generated NL prompt, extend a tag sequence.
@@
tag_to_short_to_long From tags or NL prompts, generate a refined NL prompt.
Table 1: Pre-sampling tasks in TIPO. Basic tasks focus on one-step transformations, while composite tasks combine two basic tasks within a single forward pass.
@@
We randomly select from the aforementioned tasks during training to enhance model generalization. By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences. Figure 3 (b) illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence:
[…]
@@
(a) Scenery tag only input
(b) Tag + short NL input
Figure 3: Example prompt optimization paths in TIPO. (a) shows generation from a single tag input, while (b) uses both tag and natural language input. These illustrate representative pipelines, not the full range of use cases. Blue shading indicates increasing prompt richness.
Implementation Details
In implementation, TIPO adopt the LLaMA architecture (Touvron et al., 2023a; b; AI@Meta, 2024) [^1] The multi-task design of TIPO is compatible with many autoregressive language models (e.g., GPT, LLaMA, Qwen, etc.). Intuitively, adopting more advanced backbones could further enhance efficiency and effectiveness, but such exploration would require extra training cost, which we defer to future work., with all experiments are conducted with a 200M-parameter model  [^2] We also trained TIPO-100M and 500M variants to analyze the impact of model scales. See Appendix C.. Our training dataset is about 40 billion tokens curated from Danbooru2023 (nyanko202, 2023; Yeh, 2024a), GBC10M (Hsieh et al., 2024), and CoyoHD11M (CaptionEmporium, 2024).